OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 2(n - 3) + 5 = 3(n - 1)

OpenStudy (anonymous):

2n-6+5=3n-3= =2n-1=3n-3= =2n-3n=-3+1= =-n=-2 \multiply both sides by (-1) and you'll get =n=2